ISSN 1201-561X

[figure: T; A; C]

THEORY AND APPLICATIONS OF CATEGORIES

Volume 22, 2009

Duality for CCD lattices Francisco Marmolejo, Robert Rosebrugh, and R.J. Wood 1

On deformations of pasting diagrams D. N. Yetter 24

Searching for more absolute CR-epic spaces Michael Barr, John F. Kennison, and R. Raphael 54

On endomorphism algebras of separable monoidal functors Brian Day and Craig Pastro 77

Categories with slicing Thorsten Palm 97

Tensor-triangulated categories and dualities Baptiste Calms and Jens Hornbostel 136

Lois distributives. Applications aux automates stochastiques Elisabeth Burroni 198

Algebraic colimit calculations in homotopy theory using fibred and cofibred categories Ronald Brown and Rafael Sivera 222

Covering space theory for directed topology Eric Goubault, Emmanuel Haucourt, Sanjeevi Krishnan 252 On the categorical semantics of elementary linear logic Olivier Laurent 269

\begin{matrix}c \\ ◯ \end{matrix} The Editors of Theory and Applications of Categories 2009. Permission to copy for private use

granted.

Enriched orthogonality and equivalences M. Golasinski and L. Stramaccia 302

Weak distributive laws Ross Street 313

An embedding theorem for Hilbert categories Chris Heunen 321

Eventually cyclic spectra of parameterized flows John F. Kennison 345

T-catgories reprsentables Jacques Penon 376

Epimorphic covers make R_{G}^{+} a site, for profinite G Daniel G. Davis 386

Isbell duality for modules Michael Barr, John F. Kennison, and R. Raphael 401

Higher-dimensional categories with finite derivation type Yves Guiraud and Philippe Malbos 420

A brief review of abelian categorifications Mikhail Khovanov, Volodymyr Mazorchuk and Catharina Stroppel 479

Algebraic categories whose projectives are explicitly free Matas Menni 509

Vector fields and flows on differentiable stacks Richard Hepworth 542

Homotopical interpretation of globular complex by multipointed d-space Philippe Gaucher 588

Cartesian differential categories R.F. Blute, J.R.B. Cockett and R.A.G. Seely 622

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

Full text of the journal is freely available in .dvi, Postscript and PDF from the journal’s server at http://www.tac.mta.ca/tac/ and by ftp. It is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive by e-mail abstracts of articles by e-mail as they are published. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX 2ε is strongly encouraged. Articles should be submitted by e-mail directly to a Transmitting Editor. Please obtain detailed information on submission format and style files from http://www.tac.mta.ca/tac/.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@math.mcgill.ca

Assistant TEX editor. Gavin Seal, McGill University: gavin seal@fastmail.fm

Transmitting editors.

Clemens Berger, Université de Nice-Sophia Antipolis, cberger@math.unice.fr Richard Blute, Université d’ Ottawa: rblute@uottawa.ca Lawrence Breen, Université de Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: ronnie.profbrown (at) btinternet.com Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Cuill Inc.: valeria@cuill.com Ezra Getzler, Northwestern University: getzler(at)northwestern(dot)edu Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Tom Leinster, University of Glasgow, T.Leinster@maths.gla.ac.uk Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz Brooke Shipley, University of Illinois at Chicago: bshipley@math.uic.edu James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca